OpenStudy (anonymous):

Find two integers between -10 and 10 that make the equation false

OpenStudy (paxpolaris):

what equation???

OpenStudy (anonymous):

it doesn't say thats all it says

OpenStudy (anonymous):

oh theres a part a that says for the following equation, find two integers between -10 and 10 that make the equation true: -2x^2-2x=-4

OpenStudy (paxpolaris):

\[\Large -2x^2-2x=-4\] \[\implies \Large2x^2+2x-4=0\]\[\implies \Large2\left( x^2+x-2 \right)=0\]

OpenStudy (paxpolaris):

Factor the Left side to get two number for which the equation is true. any other numbers ... the equation will be false

OpenStudy (anonymous):

what about when it says find two integers between -10 and 10 that make the equation false

OpenStudy (paxpolaris):

first you have to find for which is it true.

OpenStudy (paxpolaris):

whic are those?

OpenStudy (anonymous):

i don't know hot to find them.

OpenStudy (paxpolaris):

can you factor: \[ \Large x^2+x-2 \]

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (paxpolaris):

we had: \[\Large2\left( x^2+x-2 \right)=0\] divide both sides by 2:\[\implies \Large x^2+x-2 =0\] \[\implies \Large x^2-x\ +\ 2x-2 =0\]\[\implies \Large x \left( x-1 \right)+2\left( x-1 \right) =0\] \[\implies \huge (x+2)(x-1) =0\]

OpenStudy (paxpolaris):

so, x+2 = 0 , or x-1 =0 \(\huge x=-2\), or \(\huge x=1\) plug either of these in the original equation: \(-2x^2-2x=-4\) and you will get \(\large -4=-4\) take any other value for x (like x=0, 3, or -1) an the equation will be false